In a certain factory, three machines A, B and C produce the same product. The machines A and B produce respectively 35 and 45 percent of the total output. The percentage of defectives produced by each of the machines A, B and C are respectively 5, 4 and 2 percent.
(a) An item is drawn at random from the total output. What is the probability that it is defective?
(b) An item drawn at random from the total output turned out to be non defective. What is the probability that it was manufactured by machine A?

Answer by richwmiller(17219)

300
.35,.45,.20
.05,.04,.02
.35*300=105
.45*300=135
.20*300= 60
.05*105=5.25
.04*135=5.40
.02*60=1.2
11.85
5.25+5.40+1.2=11.85
11.85/300=0.0395
a) 3.95%
.95*105=99.75
.96*300=95.6
.98*300=98.8
294.15
99.75/294.15=0.339
b) 3.39%